# Unveiling the Role of ChatGPT in Software Development: Insights from Developer-ChatGPT Interactions on GitHub

### v1

@@ 3.3. Data Cleaning
After collecting accessible data from GitHub, we performed preliminary data cleaning to eliminate invalid search results based on the following Exclusion (E) criteria. Statistical results for the exclusions are presented in the Invalid URLs and Non-English columns of Table 1.
E1. Invalid URLs: Those invalid ChatGPT conversation URLs could be retrieved in the initial collected data for various reasons: 1) Canceled Sharing: the users who shared the conversation records may have subsequently canceled the sharing, rendering the URLs inaccessible; 2) Processed URLs: the retrieval results may include processed ChatGPT conversation URLs, resulting in entries like “chat.openai.com/share/xxx [invalid URL hashcode]”, which do not correspond to valid conversation links; 3) Invalid URL formant: some users shared ChatGPT conversation links with incorrect formats, such as “chat.openai.com/share/e/5ec5788d-8151-4c7e-ab1e-68d84a80b79f”, in which the inclusion of “e/” deviates from the standard format. Even after removing “e/”, the links remained inaccessible. Due to link disruption, we treated those instances as invalid links, as their original content could not be retrieved.
E2. Non-English text: The search results contain non-English text (except emoji), shown in the Non-English column of Table 1. Developers might use ChatGPT for tasks such as document internationalization; therefore, we excluded these non-English search results.
@@ 3.4. Data Labeling
To streamline subsequent data labeling, we extracted relevant fields (see Section 3.5) from the JSON files obtained in the previous step. These fields were primarily selected to facilitate the data labeling process and were stored in MS Excel sheets for annotation.
After the initial data cleaning, we further manually checked and labeled the initial cleaned data during the data labeling process. Meanwhile, we also needed to remove duplicate data. For example, in many cases, developers tend to create or modify multiple files within a single commit. However, the corresponding commit messages often only provide a single shared ChatGPT link, leading to multiple GitHub links for different files being associated with the same shared developer-ChatGPT conversation. To address this, it is necessary to retain only one data entity related to the shared ChatGPT links and remove redundant data entities. Our approach to data preprocessing, specifically the removal of duplicate data entities, slightly differs from that of other studies (e.g., studies based on the DevGPT dataset (Xiao et al., 2024)). To achieve our goal of understanding how developers utilize ChatGPT during the development process, we eliminated duplicate ChatGPT links (data entities sharing a single ChatGPT URL across different GitHub links). This ensures the robustness of our statistical results and data analysis.
@@ 3.5. Data Analysis
Our DevChat dataset was curated from five sources on GitHub, as detailed in Table 1. For each data source, we provided distinct metadata in JSON files to support the data analysis. Each JSON file comprises data entities with the common data items, including basic GitHub repository information (e.g., Type, URL, Author, Title, RepoName), timestamps (e.g., AuthorAt, CreatedAt), and the extracted ChatGPT data item (i.e., ChatgptSharing). Note that, ChatgptSharing consists of detailed information related to shared ChatGPT links, such as URL of shared ChatGPT conversations, timestamps of the URLs (i.e., DateOfConversation, DateOfAccess), titles generated by ChatGPT (i.e., Title), detailed statistics of the prompts and answers (i.e., NumberOfPrompts, TokenOfPrompts, TokenOfAnswers), and the model version of ChatGPT used (i.e., Model).
@@
In this work, we examine how developers leverage ChatGPT throughout the software development process. For the data analysis, we utilized statistical methods to address the characteristics (RQ1) associated with shared ChatGPT links. To examine and classify the purposes (RQ2), activities (RQ3), and SE tasks (RQ4) of employing ChatGPT on software development, we adopted the Constant Comparison method (Charmaz, 2014; Stol et al., 2016). According to Charmaz et al. (Charmaz, 2014), the Constant Comparison process involves three key steps. In the first step, initial coding, three authors (the first, third, and fourth authors) independently reviewed the content of shared ChatGPT conversations. This was followed by focused coding, where the three authors selected the most frequent codes and categorized the data accordingly, with the results subsequently reviewed by the first author. For instance, we consolidated codes such as “Dataset Split”, “Data Merge”, and “Data Cleaning” into “Data Handling”, and merged “Program Comprehension” into “Code Explanation and Understanding”. Third, we conducted theoretical coding to establish relationships between the identified codes. To minimize personal bias, disagreements in coding results were resolved through collaborative, multi-turn discussions among all the authors, ensuring an objective interpretation of the data. Note that, we extracted and stored relevant data entities from shared ChatGPT links in MS Excel files (see the dataset at (Li et al., 2025)). Our dataset also lends itself to further analysis of other topics, such as prompts, responses, and relevant interactions.
@@ 4.1. RQ1: Characteristics of ChatGPT Usages
Notably, Code emerges as the dominant share of interactions, comprising 43.4% (1,105) of all shared URLs. This significant share underscores ChatGPT’s primary role in supporting code-related tasks, such as generating code snippets or troubleshooting issues within the codebase. Commits follow as the second largest share at 32.3% (823), indicating that developers also turn to ChatGPT for support in committing changes, potentially refining, validating, or documenting code changes before integration. The remaining sources show comparatively lower levels of ChatGPT engagement. Issues account for 11.6% (295), implying moderate use of ChatGPT for tracking and resolving problems. Pull Requests contribute 10.4% (266), suggesting that developers sometimes employ ChatGPT during code review and merging processes. Lastly, Discussions make up a mere 2.3% (58), reflecting limited reliance on ChatGPT for open-ended or collaborative discussions. The distribution of shared ChatGPT links reveals a strong preference for ChatGPT in focused and procedural work like coding and committing. In contrast, its use in more interactive and exploratory exchanges is comparatively lower, such as discussions.
@@
From another perspective, Figure 3 illustrates the cumulative adoption of ChatGPT by developers. The steady growth of ChatGPT’s usage across the five sources highlights the platform’s popularity and importance in the software development community. The lines for Code and Commits consistently remain the most prominent, which aligns with GitHub’s primary role as a code repository and a platform for version control. The lines for Issues and Pull Requests show a steady increase, reflecting that collaborative development and knowledge sharing are prevalent on GitHub. The Discussions line remains relatively low compared to the others, suggesting that discussions might not be as heavily utilized on GitHub as other activities.
Overall, this trend reflects ChatGPT’s emerging role in software development, especially in collaborative and code-centric work on GitHub.
@@
Our DevChat dataset also contains the prompt data of developer-ChatGPT interactions. Figure 4 provides a box plot visualization of the prompt turns distribution across five GitHub sources, after filtering outliers. The associated statistics are summarized in Table 2, which includes the first (Q1), second (Q2, a.k.a Median), and third (Q3) quartiles of the distributions.
The high average number of prompt turns in shared ChatGPT conversations observed in Code indicates the inherently iterative and complex nature of missions like code generation, debugging, and explanation. Moreover, the prompt turns of developer-ChatGPT interactions in Code are higher than those found in other data sources. The distribution of prompt turns in Commits, Issues, Pull Requests, and Discussions indicates developers’ usage of ChatGPT for short and task-focused interactions (e.g., concept interpretation, code modification).
@@
Source | Count | Q1 | Q2 | Q3 | Min/Max |
Code | 1,105 | 1 | 3 | 7 | 1 / 302 |
Commits | 823 | 1 | 1 | 2 | 1 / 89 |
Issues | 295 | 1 | 2 | 4 | 1 / 109 |
Pull Requests | 266 | 1 | 2 | 4 | 1 / 167 |
Discussions | 58 | 1 | 2 | 4 | 1 / 82 |
@@
To investigate the descriptions of shared ChatGPT links, we examined the description distribution of the shared ChatGPT links across the five GitHub sources. We identified two primary categories when developers share ChatGPT links: sharing ChatGPT links with or without a description. Figure 5 presents an overview of the two categories regarding developer-ChatGPT interactions.
Figure 5 reveals a strong tendency to include contextual description when sharing ChatGPT links, especially in collaborations like Commits (99.39%), Discussions (91.38%), and Pull Requests (89.72%). In contrast, the practices of attaching descriptions to shared ChatGPT links in Code vary, 70.95% and 29.05% of the shared ChatGPT links with and without a description, respectively. More specifically, nearly half (47.33%) of the shared ChatGPT links in code comments provide a contextual description, while 23.17% lack contextual information. Shared links lacking descriptions are minimal in README files (2.62%) and code snippets (0.81%). This indicates that developers may rely on the link content to interpret or assume that the recipient has the necessary background to understand the link content.
Overall, 82.80% of shared ChatGPT links contain descriptions, whereas 17.20% of them do not offer any contextual information. The findings reflect that developers generally prefer to provide descriptive information when sharing ChatGPT links, possibly to ensure clarity and facilitate collaborations in development with the shared developer-ChatGPT interactions.
@@ 4.2. RQ2: Developers’ Purposes for Sharing ChatGPT Conversations
To answer RQ2, we employed the Constant Comparison method (Charmaz, 2014; Stol et al., 2016) to analyze and categorize the developers’ purposes for sharing ChatGPT links during software development. We finally got five primary categories of developers’ purposes (i.e., Task Delegation, Problem Resolution, Knowledge Acquisition, Solution Recommendation, and Concept Interpretation), and we analyzed their prevalence across the five sources. Figure 6 presents the distribution of the five purposes across five sources. Each circle is scaled to the total number of shared ChatGPT links for different sources, and the segments within each circle show the relative contribution of each purpose, making cross-category comparisons intuitive. Detailed results are presented below.
@@
(1) Task Delegation encompasses the use of ChatGPT to automate or execute specific development tasks. Task delegation emerges as the predominant purpose, particularly in Commits (85.66%) and Pull Requests (88.35%), reflecting ChatGPT’s role in automating repetitive tasks. For instance, developers frequently employ ChatGPT to generate boilerplate code and README files. The high prevalence of task delegation in Code (75.02%) further underscores the utility of task delegation in accelerating code creation. However, the lower adoption of task delegation in Issues (61.02%) and Discussions (50.00%) suggests that collaborative or context-rich tasks may demand human inspection, limiting the automation potential in these sources.
@@
(2) Problem Resolution involves debugging code, fixing errors, or explaining error messages. Problem resolution peaks in Commits (11.42%) and Issues (8.47%), aligning with scenarios where developers address problems when merging code or triaging issues. The minimal presence of problem resolution in Pull Requests (1.13%) implies that ChatGPT may not be as commonly relied upon for resolving complex or collaborative challenges. However, ChatGPT’s utility in problem resolution underscores its potential as an efficient tool to identify and fix errors during development.
@@
(3) Knowledge Acquisition refers to the proactive process that developers seek for clarifications, differences, and domain-specific knowledge during development to contribute effectively to their projects. Knowledge acquisition is relatively prominent in Discussions (20.69%) and Issues (12.88%), where developers seek clarifications or compare technical approaches relevant to development. While ChatGPT’s role in knowledge acquisition is not as dominant in Commits (0.61%) and Pull Requests (4.51%), ChatGPT’s contribution to the educational aspects of development is noteworthy. By serving as a knowledge repository and a search engine, ChatGPT helps developers quickly access relevant details and broaden their understanding of various knowledge points.
@@
(4) Solution Recommendation covers requests for solutions, recommendations, suggestions, etc, to guide implementation or resolve issues efficiently. Solution Recommendation is comparatively more frequent in Discussions (17.24%) and Issues (12.54%), indicating ChatGPT’s value in collaborative problem-solving. Developers appear to use ChatGPT to generate actionable suggestions during ideation, such as recommending design patterns, best practices, and decision-making for making sound technical decisions. The lower prevalence of this task in Code (2.26%) and Pull Requests (2.63%) implies that GitHub’s usage conventions that developers focus on execution (e.g., implement predefined and specific tasks like bug fixing) rather than exploration (e.g., recommend and try new possibilities like designing algorithms) in these two sources on GitHub (Wessel et al., 2023).
@@
(5) Concept Interpretation: Developers usually seek ChatGPT’s help to understand or interpret specific concepts, algorithms, design principles, and other software development topics. Concept interpretation is the least adopted category overall, with modest usage in Issues (5.08%) and Discussions (5.17%). These requests typically highlight ChatGPT’s role as a development assistant. It reflects ChatGPT’s strength in facilitating concept comprehension, making ChatGPT an invaluable tool for developers seeking quick explanations or insights on unfamiliar concepts.

### v2

@@ 3.3. Data Cleaning
After collecting accessible data from GitHub, we performed preliminary data cleaning to eliminate invalid search results based on the following Exclusion (E) criteria. The results for the exclusions are presented in the Invalid URLs and Non-English columns of Table 2.
E1. Invalid URLs: Those invalid shared ChatGPT links could be retrieved in the initial collected data for various reasons: 1) Canceled Sharing: the users who shared the conversation records may have subsequently canceled the sharing, rendering the URLs inaccessible; 2) Processed URLs: the retrieval results may include processed shared ChatGPT links, resulting in entries like “chat.openai.com/share/xxx [invalid URL hashcode]”, which do not correspond to valid conversation links; 3) Invalid URL formant: some users shared ChatGPT conversation links with incorrect formats, such as “chat.openai.com/share/e/5ec5788d-8151-4c7e-ab1e-68d84a80b79f”, in which the inclusion of “e/” deviates from the standard format. Even after removing “e/”, the links remained inaccessible. Due to link disruption, we treated those instances as invalid links, as their original content could not be retrieved.
E2. Non-English text: The conversations in the retrieved shared ChatGPT links contain non-English text (except emoji), shown in the Non-English column of Table 2. Developers might use ChatGPT for tasks such as document internationalization; therefore, we excluded these non-English search results. Note that code snippets were not considered in data cleaning, meaning that we did not remove any shared ChatGPT links by their contained code snippets during data cleaning.
Duplicates Removal. After the abovementioned data cleaning by applying the Exclusion criteria E1 and E2, we manually inspected the dataset and removed duplicate data. For example, developers often add or modify multiple files within a single commit; when ChatGPT links are pasted into code comments, the same ChatGPT URL can appear across several GitHub file links, as shown in Figure 2. Meanwhile, we have manually checked and removed the ChatGPT links that are not relevant to the content of the data sources where the links are located. Moreover, to avoid counting the same conversation multiple times, we retained only one record per unique ChatGPT link and removed the redundant records. We have conducted a deduplication procedure to eliminate records that share a single ChatGPT URL across different GitHub links, and the deduplication enhances the robustness of our results and data analysis.
@@ 3.4. Data Labeling
To streamline subsequent data labeling, extraction, and analysis, we examined developer-ChatGPT conversations in the cleaned dataset, checking the shared ChatGPT links and their contextual information within GitHub repositories. The data labeling process aims to identify and label the developer-ChatGPT interactions that are directly relevant to software development.
@@ 3.5. Data Extraction and Analysis
Our DevChat dataset was curated from five sources on GitHub, as detailed in Table 2. For each data source, we provided distinct metadata in JSON files to support the data analysis. Each JSON file comprises data entities with the common data items, including (1) basic GitHub repository information (for answering RQ1) (e.g., Source Type, URL, Author, Title, RepoName), timestamps (e.g., AuthorAt, CreatedAt), and (2) the extracted ChatGPT data items (for answering RQ2, RQ3, RQ4) (i.e., ChatgptSharing). Note that, ChatgptSharing consists of detailed information related to shared ChatGPT links, such as URLs of shared ChatGPT conversations, timestamps of the URLs (i.e., DateOfConversation, DateOfAccess), titles of ChatGPT conversations (i.e., Title) and corresponding content, detailed statistics of the prompts and answers (i.e., NumberOfPrompts, TokenOfPrompts, TokenOfAnswers), and the model version of ChatGPT used (i.e., Model).
@@
In this work, we examined how developers leverage ChatGPT throughout the software development process. For answering RQ1, we presented the distributions and characteristics of the shared ChatGPT links on GitHub, and we further analyzed the contents of shared ChatGPT conversations to answer the rest RQs (i.e., RQ2, RQ3, and RQ4). For the quantitative analysis addressing characteristics (RQ1) of shared ChatGPT links, we applied descriptive statistics. Concretely, we computed frequencies and percentages of the shared ChatGPT links, calculated the central tendency (mean, median, and mode) and dispersion (including the interquartile range), and visualized results distributions with histograms and bar plots. To examine and classify developers’ purposes (RQ2), activities (RQ3), and tasks (RQ4) when sharing ChatGPT links, we adopted the Constant Comparison method (Charmaz, 2014; Stol et al., 2016), a grounded theory technique rigorously validated for deriving emergent categories from qualitative data in software engineering research. Constant Comparison was chosen specifically because it is uniquely suited for exploratory qualitative research where categories (e.g., developers’ purposes) must emerge inductively from unstructured data, rather than being imposed a priori. According to the description provided in (Charmaz, 2014), the Constant Comparison process consists of three main steps: (1) initial coding, three authors (the first, third, and fourth authors) independently reviewed the content of shared ChatGPT conversations. For example, we initially labeled 22 purposes for RQ2 before the subsequent label merge (i.e., focused coding). (2) focused coding, the three authors selected the most frequent codes and categorized the data accordingly, with the results subsequently reviewed by the first author. For instance, we consolidated codes such as “Dataset Split”, “Data Merge”, and “Data Cleaning” into “Data Handling”, and merged “Program Comprehension” into “Code Explanation and Understanding”. (3) theoretical coding, the three authors aimed to establish relationships between the identified codes. For instance, we grouped the “Lifecycle Activities” and “Supporting Activities” into the root node “Development Related Activities” for RQ3. To minimize personal bias, disagreements in coding results were resolved through collaborative, multi-turn discussions among all the authors, ensuring an objective interpretation of the data. Note that, we extracted and stored relevant data entities from shared ChatGPT links in MS Excel files (see the dataset at (Li et al., 2025)) to facilitate data analysis. Our dataset also lends itself to further analysis of other topics, such as prompts, responses, and relevant interactions.
@@ 4.1. RQ1: Characteristics of ChatGPT Usages
Notably, Code emerges as the dominant share of interactions, comprising 43.4% (1,105) of all shared URLs. This reflects that developers most frequently use ChatGPT in contexts directly related to source code, such as generating or modifying code snippets. Commits follow as the second largest share at 32.3% (823), indicating that shared ChatGPT links are commonly referenced during the commit process. The remaining sources show comparatively lower usage: Issues account for 11.6% (295), Pull Requests for 10.4% (266), Discussions for only 2.3% (58), reflecting limited reliance on ChatGPT for open-ended or collaborative discussions. Overall, the distribution indicates that developer-ChatGPT conversations are most frequently shared in contexts directly related to code development and integration activities, while its presence in collaborative or discussion-oriented contexts is comparatively lower.
@@
From another perspective, Figure 4 illustrates the cumulative usage of ChatGPT by developers. Specifically, the vertical axis in Figure 4 shows the cumulative usage of ChatGPT from each source every month (see Figure 3). It not only highlights the monthly growth trend but also reflects the total cumulative usage as of a specific month, which provides a more detailed view of the comparison among the five data sources. The steady growth of ChatGPT usage across the five sources highlights the platform’s popularity and importance in the software development community. The lines for Code and Commits consistently remain the most prominent, which aligns with GitHub’s primary role as a code repository and a platform for version control. The lines for Issues and Pull Requests show a steady increase, reflecting that collaborative development and knowledge sharing are prevalent on GitHub. The Discussions line remains relatively low compared to the others, suggesting that discussions might not be as heavily utilized on GitHub as other activities.
Overall, Figure 3 and Figure 4 provide complementary perspectives, enabling a better understanding of the data distribution characteristics from different visual angles. This distribution reflects ChatGPT’s emerging role in software development, especially in collaborative and code-centric work on GitHub.
@@
Our DevChat dataset also contains the prompt data of developer-ChatGPT interactions. In this context, a turn refers to one complete prompt-response pair between a developer and ChatGPT (i.e., one developer’s question followed by one ChatGPT’s response). Figure 5 provides a box plot visualization of the prompt turns distribution across five GitHub sources, after filtering outliers. The corresponding statistics are summarized in Table 3, which includes the first (Q1), second (Q2, a.k.a Median), third (Q3) quartiles of the distributions, and the InterQuartile Range (IQR), along with the mode and its corresponding frequency. Together, Figure 5 enables a rapid visual assessment of the prompt turns distribution, and Table 3 provides a precise quantitative comparison of the prompt turn numbers.
The median number (Q2) of prompt turns in Code interactions (3) exceeds those in other GitHub sources, with a significantly wider interquartile range (IQR = 6). The median number indicates that code-related tasks (e.g., generation, debugging, and explanation) typically involve iterative, multi-turn dialogues requiring sustained collaboration between developers and ChatGPT. Conversely, the low median values and narrow IQRs observed in Commits, Issues, Pull Requests, and Discussions reflect predominantly short and task-focused developer-ChatGPT interactions (e.g., concept interpretation, code modification).
@@
| Source | Count | Q1 | Q2 | Q3 | IQR | Min/Max | Mode/Frequency |
| Code | 1,105 | 1 | 3 | 7 | 6 | 1 / 302 | 1 / 298 |
| Commits | 823 | 1 | 1 | 2 | 1 | 1 / 89 | 1 / 441 |
| Issues | 295 | 1 | 2 | 4 | 3 | 1 / 109 | 1 / 107 |
| Pull Requests | 266 | 1 | 2 | 4 | 3 | 1 / 167 | 1 / 104 |
| Discussions | 58 | 1 | 2 | 4 | 3 | 1 / 82 | 1 /25 |
@@
To investigate the descriptions of shared ChatGPT links (e.g., the description in the code comment in Figure 2), we quantified the distribution of links with versus without contextual descriptions across the five GitHub sources in Figure 6. In this study, a “description” refers to the accompanying textual explanation or context provided alongside a shared ChatGPT link, such as a comment, note, or sentence clarifying the purpose, content, or relevance of the linked ChatGPT interaction. We identified two primary categories when developers share ChatGPT links: sharing ChatGPT links with or without a description.
Figure 6 reveals a strong tendency to include contextual description when sharing ChatGPT links, especially in collaborations like Commits (99.39%), Discussions (91.38%), and Pull Requests (89.72%). In contrast, the practices of attaching descriptions to shared ChatGPT links in Code vary, 70.95% and 29.05% of the shared ChatGPT links with and without a description, respectively. Overall, 82.80% of shared ChatGPT links contain descriptions, whereas 17.20% of them do not offer any contextual information. The findings reflect that developers generally prefer to provide descriptive information when sharing ChatGPT links, possibly to ensure clarity and facilitate collaborations in development with the shared developer-ChatGPT interactions.
@@ 4.2. RQ2: Developers’ Purposes for Sharing ChatGPT Conversations
To answer RQ2, we employed the Constant Comparison method (Charmaz, 2014; Stol et al., 2016) to analyze and categorize the developers’ purposes for sharing ChatGPT links during software development. Specifically, we examined the textual context surrounding the shared links, including the link description, following messages, and explicit user intent statements; for those links without descriptions, we analyzed their developer-ChatGPT conversations. We finally got five primary categories of developers’ purposes (i.e., Task Delegation, Problem Resolution, Knowledge Acquisition, Solution Recommendation, and Concept Interpretation), and we analyzed their prevalence across the five sources. Figure 7 presents the distribution of the five purposes across five sources. Each circle is scaled to the total number of shared ChatGPT links for different sources, and the segments within each circle show the relative contribution of each purpose, making cross-category comparisons intuitive. For each category, we provide detailed descriptions of the results, followed by one example from our curated dataset. Detailed results are presented below.
@@
(1) Task Delegation: Use of ChatGPT to automate or execute specific development tasks (e.g., generating unit tests, refactoring functions). Task delegation emerges as the predominant purpose, particularly in Commits (85.66%) and Pull Requests (88.35%), reflecting ChatGPT’s role in automating repetitive tasks. For instance, developers frequently employ ChatGPT to generate boilerplate code and README files. The high prevalence of task delegation in Code (75.02%) further underscores the utility of task delegation in accelerating code creation. However, the lower adoption of task delegation in Issues (61.02%) and Discussions (50.00%) suggests that collaborative or context-rich tasks may demand human inspection, limiting the automation potential in these sources.
@@
(2) Problem Resolution: Interactions aim at diagnosing and resolving faults, including debugging, fixing runtime errors, and explaining error messages. Problem resolution peaks in Commits (11.42%) and Issues (8.47%), aligning with scenarios where developers address problems when merging code or triaging issues. The minimal presence of problem resolution in Pull Requests (1.13%) implies that ChatGPT may not be as commonly relied upon for resolving complex or collaborative challenges. However, ChatGPT’s utility in problem resolution underscores its potential as an efficient tool to identify and fix errors during development.
@@
(3) Knowledge Acquisition: Proactive requests for clarification, background information, or domain-specific knowledge that help developers reason about design or implementation choices. Knowledge acquisition is relatively prominent in Discussions (20.69%) and Issues (12.88%), where developers seek clarifications or compare technical approaches relevant to development. While ChatGPT’s role in knowledge acquisition is not as dominant in Commits (0.61%) and Pull Requests (4.51%), ChatGPT’s contribution to the educational aspects of development is noteworthy. By serving as a knowledge repository and a search engine, ChatGPT helps developers quickly access relevant details and broaden their understanding of various knowledge points.
@@
(4) Solution Recommendation: Requests for options, trade-offs, design suggestions, or concrete corrective actions to guide implementation or choice among alternatives. Solution Recommendation is comparatively more frequent in Discussions (17.24%) and Issues (12.54%), indicating ChatGPT’s value in collaborative problem-solving. Developers tend to use ChatGPT to generate actionable suggestions during ideation, such as recommending design patterns, best practices, and decision-making for making sound technical decisions. The lower prevalence of this task in Code (2.26%) and Pull Requests (2.63%) implies that GitHub’s usage conventions that developers focus on execution (e.g., implement predefined and specific tasks like bug fixing) rather than exploration (e.g., recommend and try new possibilities like designing algorithms) in the two sources (Wessel et al., 2023).
@@
(5) Concept Interpretation: Requests to explain, interpret, or translate specific technical concepts (e.g., algorithms, design patterns, protocols, or libraries) often at a conceptual level. Developers usually seek ChatGPT’s help to understand or interpret specific concepts, algorithms, design principles, and other software development topics. Concept interpretation is the least adopted category overall, with modest usage in Issues (5.08%) and Discussions (5.17%). These requests typically highlight ChatGPT’s role as a development assistant. It reflects ChatGPT’s strength in facilitating concept comprehension, making ChatGPT an invaluable tool for developers seeking quick explanations or insights on unfamiliar concepts.